# Spectral theory of the Neumann-Poincaré operator associated with multi-layer structures and analysis of plasmon mode splitting

### v2

@@ 3 Properties of $\mathbb{K}_{N}^{*}$
$\mathcal{S}_{S_{e}}\Big{[}\frac{\partial u}{\partial\nu}\Big{|}_{-}\Big{]}(x)-\mathcal{S}_{S_{i}}\Big{[}\frac{\partial u}{\partial\nu}\Big{|}_{+}\Big{]}(x)=\mathcal{D}_{S_{e}}\Big{[}u\big{|}_{-}\Big{]}(x)-\mathcal{D}_{S_{e}}\Big{[}u\big{|}_{+}\Big{]}(x),\quad x\in\mathbb{R}^{d}\setminus\overline{B}.$
@@
$-\frac{1}{2}\mathcal{S}_{\partial B}[\phi](x)+\mathcal{S}_{\partial B}\mathcal{K}_{\partial B}^{*}[\phi](x)=\mathcal{D}_{\partial B}\mathcal{S}_{\partial B}[\phi](x),\quad x\in\mathbb{R}^{d}\setminus\overline{B}.$
@@
[…]
@@
    The spectrum of $\mathbb{K}_{N}^{*}$ on $\mathcal{H}^{-\frac{1}{2}}$ lies in the interval $[-1/2,1/2]$.
@@
which show that $\mathbb{K}_{N}^{*}$ is self-adjoint on $\mathcal{H}$. Moreover, the the norm $\|\cdot\|_{\mathcal{H}}$ induced by the inner product (3.18) is equivalent to the $\mathcal{H}^{-\frac{1}{2}}_{0}$ norm. We refer to [27] for more detail. Since the interfaces $\Gamma_{k}$, $k=1,2,\ldots,N$, are of class $C^{1,\eta}$ for $10<\eta<$ 1, the NP operator $\mathbb{K}_{N}^{*}$ is compact as well as self-adjoint on $\mathcal{H}$. Hence the spectrum $\sigma(\mathbb{K}_{N}^{*})$ is real and has only point spectrum accumulating to 0.
@@
$\quad\mathcal{S}_{\Gamma_{l}}\Big{[}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}\Big{]}(x)-\mathcal{S}_{\Gamma_{n}}\Big{[}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{n}}\Big{|}_{+}\Big{]}(x)$
@@
$\quad\mathcal{S}_{\Gamma_{l}}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}-\frac{1}{2}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]-\mathcal{S}_{\Gamma_{n}}\mathcal{K}_{\Gamma_{n}}^{*}[\phi_{n}]$ (3.24)
@@
$-\frac{1}{2}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]+\mathcal{S}_{\Gamma_{n}}\mathcal{K}_{\Gamma_{n}}^{*}[\phi_{n}]=\mathcal{D}_{\Gamma_{n}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]\quad\textmd{on }\Gamma_{m},$ (3.25)
@@
$\quad\mathcal{S}_{\Gamma_{n}}\Big{[}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{n}}\Big{|}_{-}\Big{]}(x)-\mathcal{S}_{\Gamma_{l}}\Big{[}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}\Big{]}(x)$
@@
$\quad-\frac{1}{2}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]+\mathcal{S}_{\Gamma_{n}}\mathcal{K}_{\Gamma_{n}}^{*}[\phi_{n}]-\mathcal{S}_{\Gamma_{l}}\frac{\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}$ (3.30)
@@
$=(-1)^{m}\mathcal{K}_{\Gamma_{m}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]+\sum\limits_{m<l\leqslant n}(-1)^{l}\mathcal{D}_{\Gamma_{l}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]$
@@
4 Explicit formulae for the integral equation in multi-layer confocal ellipses
In this section, we explicitly compute the solution $\phi_{k}$ of the integral equation (2.8) on the non-radial case, more specifically, we consider the conductivity transmission problem (2.1)–(2.4) with $A$ being a multi-layer confocal ellipse. We introduce the elliptic coordinates $(\xi,\eta)$ so that $x=(x_{1},x_{2})$ in Cartesian coordinates are defined by
@@ 4 Explicit formulae for the integral equation in multi-layer confocal ellipses
where $N\in\mathbb{N}$. We assume $\Gamma_{k},$ $k=1,2,\ldots,N$, are confocal ellipses whose common foci are denoted by (±$(\pm R,0)$. So the confocal ellipses have the same elliptic coordinates. The illustration of $N$-layer confocal ellipses is shown in Figure 4.1.
Figure 4.1: Schematic illustration of an $N$-layer confocal ellipse.
For any positive elliptic radius $\xi_{0}$, the set $\Gamma_{0}=\{(\xi,\eta):\xi=\xi_{0}\}$ is an ellipse whose foci are (±$(\pm R,0)$. One can see easily that the length element $\mathrm{d}s$ and the outward normal derivative $\frac{\partial}{\partial\nu}$ on $\Gamma_{0}$ are given in terms of the elliptic coordinates by
@@
$\begin{split}\mathbf{M}_{N,c}^{(n)}:=\begin{bmatrix}\lambda_{1}-\frac{1}{2\mathrm{e}^{2n\xi_{1}}}&-\frac{\cosh(n\xi_{2})}{\mathrm{e}^{n\xi_{1}}}&\cdots&-\frac{\cosh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{1}}}&-\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{1}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\sinh(n\xi_{2})}{\mathrm{e}^{n\xi_{1}}}&\lambda_{2}-\frac{1}{2\mathrm{e}^{2n\xi_{2}}}&\cdots&-\frac{\cosh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{2}}}&-\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{2}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\vdots&\vdots&\ddots&\vdots&\vdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\sinh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{1}}}&\frac{\sinh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{2}}}&\cdots&\lambda_{N-1}-\frac{1}{2\mathrm{e}^{2n\xi_{N-1}}}&-\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{N-1}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{1}}}&\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{2}}}&\cdots&\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{N-1}}}&\lambda_{N}-\frac{1}{2\mathrm{e}^{2n\xi_{N}}}\end{bmatrix}.\end{split}$ (4.41)
@@
$\begin{split}\mathbf{M}_{N,s}^{(n)}:=\begin{bmatrix}\lambda_{1}+\frac{1}{2\mathrm{e}^{2n\xi_{1}}}&-\frac{\sinh(n\xi_{2})}{\mathrm{e}^{n\xi_{1}}}&\cdots&-\frac{\sinh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{1}}}&-\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{1}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\cosh(n\xi_{2})}{\mathrm{e}^{n\xi_{1}}}&\lambda_{2}+\frac{1}{2\mathrm{e}^{2n\xi_{2}}}&\cdots&-\frac{\sinh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{2}}}&-\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{2}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\vdots&\vdots&\ddots&\vdots&\vdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\cosh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{1}}}&\frac{\cosh(n\xi_{N-1})}{\mathrm{e}^{n\xi_{2}}}&\cdots&\lambda_{N-1}+\frac{1}{2\mathrm{e}^{2n\xi_{N-1}}}&-\frac{\sinh(n\xi_{N})}{\mathrm{e}^{n\xi_{N-1}}}\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{1}}}&\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{2}}}&\cdots&\frac{\cosh(n\xi_{N})}{\mathrm{e}^{n\xi_{N-1}}}&\lambda_{N}+\frac{1}{2\mathrm{e}^{2n\xi_{N}}}\end{bmatrix}.\end{split}$ (4.42)
@@
Let the multi-layer confocal ellipses $A=\cup_{k=1}^{N}A_{k}$ be given by (4.34)–(4.35). Assume that the background electrical potential $H$ admits the following multipolar expansion:
@@
[…]
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with $A$ being the multi-layer confocal ellipses given by (4.34)–(4.35). Let $H$ be given by (4.43). If $\mathbf{M}_{N,c}^{(n)}$ is invertible, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
$u-H=-\sum_{n=1}^{+\infty}a_{n,c}\cos(n\eta)\left(\cosh(n\xi)\bm{e}_{1:l}^{T}+\frac{1}{\mathrm{e}^{n\xi}}\bm{e}_{l+1:N}^{T}\right)\mathbb{F}^{(n)}_{N,l,c}\left(\mathbf{M}^{(n)}_{N,c}\right)^{-1}\bm{s},\quad\mbox{in }A_{l},\;l=0,1,\ldots,N,$
@@
It follows from Theorem 4.1, (2.7) and (4.36) that the perturbed electric potential $u-H$ in $A_{l}$, $l=0,1,\ldots,N$, can be given by
@@
Let the multi-layer confocal ellipses $A=\cup_{k=1}^{N}A_{k}$ be given by (4.34)–(4.35). Assume that the background electrical potential $H$ admits the following multipolar expansion:
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with $A$ being the multi-layer confocal ellipses given by (4.34)–(4.35). Let $H$ be given by (4.45). If $\mathbf{M}_{N,s}^{(n)}$ is invertible, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
$u-H=-\sum_{n=1}^{+\infty}a_{n,s}\sin(n\eta)\left(\sinh(n\xi)\bm{e}_{1:l}^{T}+\frac{1}{\mathrm{e}^{n\xi}}\bm{e}_{l+1:N}^{T}\right)\mathbb{F}^{(n)}_{N,l,s}\left(\mathbf{M}^{(n)}_{N,s}\right)^{-1}\bm{c},\quad\mbox{in }A_{l},\;l=0,1,\ldots,N,$
@@
One can see that the problem (2.3) is solvable if and only if the corresponding $n$-order GPM $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ is invertible. If $\Re\sigma>0$, then $\nabla\cdot\sigma(x)\nabla$ is an elliptic PDO (Partial Differential Operator), which guarantees the well-posedness of (2.3) and hence the invertibility of $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ On the other hand, if for some layers the materials are negative, the ellipticity of $\nabla\cdot\sigma(x)\nabla$ might be broken. Nevertheless, if for those negative materials, say $\sigma_{k}$, one has that $\Im\sigma_{k}\neq 0$, then the invertibility of $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ can also be guaranteed. This can be seen in our subsequent analysis.
@@ 5.1 Plasmon resonance
As mentioned above, plasmon resonance is usually associated with some eigenvalue problem generated by the PDE system. According to our earlier discussion in Remark 4.1, if the parameters $\sigma_{k}$, $k=1,2,\ldots,N$, are all positive real-valued, the elastostatic system the conductivity transmission problem (2.1)–(2.4) has only trivial solution if ${H}=0$. We seek non-trivial solutions to (2.1)–(2.4) when the conductivity in (2.10) is allowed to be negative valued, i.e.,
@@
By using Theorems 4.2 and 4.4, we can oatain the exact matrix representation of the NP operator in multi-layer confocal ellipses.
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with the conductivity $\sigma$ given by (2.10) and $A$ being the multi-layer confocal ellipses given by (4.34)–(4.35).
@@
    If $-\lambda$ is not the eigenvalue of $\mathbf{K}_{N,c}^{(n)}$, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
@@
    If $-\lambda$ is not the eigenvalue of $\mathbf{K}_{N,s}^{(n)}$, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
@@
where $-\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ or $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, then the plasmon resonance mode occurs.
@@
[…]
@@
where $\bm{\phi}=\left(\phi^{n}_{1},\phi^{n}_{2},\ldots,\phi^{n}_{N}\right)$. If $-\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$, by letting $\delta\rightarrow 0$, one has $\left|-\lambda^{*}\mathbf{I}_{N}-\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}\right|=0$ and thus the homogeneous linear equation (5.52) has non-trivial solutions. In view of $u(x)=\sum_{k=1}^{N}\mathcal{S}_{\Gamma_{k}}[\phi_{k}](x)$ and (4.44), one immediately has the problem (5.51) has non-trivial solutions.
If $-\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, one also has the problem (5.51) has non-trivial solutions in a similar manner. The proof is complete. ∎
@@
$\lim_{\delta\rightarrow 0}||\nabla({u}-{H})||_{L^{2}(\mathbb{R}^{d}\setminus\overline{A})}=+\infty.$

### v3

@@ 3 Properties of $\mathbb{K}_{N}^{*}$
$\mathcal{S}_{S_{e}}\Big[\frac{\partial u}{\partial\nu}\Big|_{-}\Big](x)-\mathcal{S}_{S_{i}}\Big[\frac{\partial u}{\partial\nu}\Big|_{+}\Big](x)=\mathcal{D}_{S_{e}}\Big[u\big|_{-}\Big](x)-\mathcal{D}_{S_{i}}\Big[u\big|_{+}\Big](x),\quad x\in\mathbb{R}^{d}\setminus\overline{B}.$
@@
$\frac{1}{2}\mathcal{S}_{\partial B}[\phi](x)+\mathcal{S}_{\partial B}\mathcal{K}_{\partial B}^{*}[\phi](x)=\mathcal{D}_{\partial B}\mathcal{S}_{\partial B}[\phi](x),\quad x\in\mathbb{R}^{d}\setminus\overline{B}.$
@@
where $\langle\cdot,\cdot\rangle$ is the pair between $\mathcal{H}^{-\frac{1}{2}}$ and $\mathcal{H}^{\frac{1}{2}}$ pairing. Then we have
@@
    The spectrum of $\mathbb{K}_{N}^{*}$ on $\mathcal{H}^{-\frac{1}{2}}$ lies in the interval [1/2,1/2][-1/2,1/2].
@@
which show that $\mathbb{K}_{N}^{*}$ is self-adjoint on $\mathcal{H}$. Moreover, the the norm $\|\cdot\|_{\mathcal{H}}$ induced by the inner product (3.4) is equivalent to the $\mathcal{H}^{-\frac{1}{2}}_{0}$ norm. We refer to [27] for more detail. Since the interfaces $\Gamma_{k}$, $k=1,2,\ldots,N$, are of class $C^{1,\eta}$ for $0<\eta<1$, the NP operator $\mathbb{K}_{N}^{*}$ is compact as well as self-adjoint on $\mathcal{H}$. Hence the spectrum $\sigma(\mathbb{K}_{N}^{*})$ is real and has only point spectrum accumulating to 0.
@@
$\quad\mathcal{S}_{\Gamma_{l}}\Big[\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}\Big](x)-\mathcal{S}_{\Gamma_{n}}\Big[\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{n}}\Big|_{+}\Big](x)$
@@
$\quad\mathcal{S}_{\Gamma_{l}}\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}-\frac{1}{2}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]-\mathcal{S}_{\Gamma_{n}}\mathcal{K}_{\Gamma_{n}}^{*}[\phi_{n}]$ (3.10)
@@
[…]
@@
$\mathcal{S}_{\Gamma_{l}}\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}=\mathcal{D}_{\Gamma_{l}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]\quad\textmd{on }\Gamma_{m}.$ (3.12)
@@
$\quad\mathcal{S}_{\Gamma_{n}}\Big[\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{n}}\Big|_{-}\Big](x)-\mathcal{S}_{\Gamma_{l}}\Big[\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}\Big](x)$
@@
$\quad-\frac{1}{2}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]+\mathcal{S}_{\Gamma_{n}}\mathcal{K}_{\Gamma_{n}}^{*}[\phi_{n}]-\mathcal{S}_{\Gamma_{l}}\frac{\partial\mathcal{S}_{\Gamma_{n}}[\phi_{n}]}{\partial\nu_{l}}$ (3.16)
@@
$=(-1)^{m}\mathcal{K}_{\Gamma_{m}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]+\sum\limits_{l\neq m}(-1)^{l}\mathcal{D}_{\Gamma_{l}}\mathcal{S}_{\Gamma_{n}}[\phi_{n}]$
@@
4 Explicit formulae for the integral equation in multi-layered confocal ellipses
In this section, we explicitly compute the solution $\phi_{k}$ of the integral equation (2.8) on the non-radial case, more specifically, we consider the conductivity transmission problem (2.1)–(2.4) with $A$ being a multi-layered confocal ellipse. We introduce the elliptic coordinates $(\xi,\eta)$ so that $x=(x_{1},x_{2})$ in Cartesian coordinates are defined by
@@ 4 Explicit formulae for the integral equation in multi-layered confocal ellipses
where $N\in\mathbb{N}$. We assume $\Gamma_{k},$ $k=1,2,\ldots,N$, are confocal ellipses whose common foci are denoted by $(\pm R,0)$. So the confocal ellipses have the same elliptic coordinates. The illustration of $N$-layer confocal ellipses is shown in Figure 4.1.
Figure 4.1: Schematic illustration of an NN-layer confocal ellipse.
For any positive elliptic radius $\xi_{0}$, the set $\Gamma_{0}=\{(\xi,\eta):\xi=\xi_{0}\}$ is an ellipse whose foci are $(\pm R,0)$. One can see easily that the length element $\mathrm{d}s$ and the outward normal derivative $\frac{\partial}{\partial\nu}$ on $\Gamma_{0}$ are given in terms of the elliptic coordinates by
@@
[…]
@@
[…]
@@
Let the multi-layered confocal ellipses $A=\cup_{k=1}^{N}A_{k}$ be given by (4.2)–(4.3). Assume that the background electrical potential $H$ admits the following multipolar expansion:
@@
$\left\{\begin{aligned} &\left(\lambda_{1}-\frac{1}{2\mathrm{e}^{2n\xi_{1}}}\right)\phi^{n}_{1}-\frac{1}{\mathrm{e}^{n\xi_{1}}}\sum_{k=2}^{N}\phi^{n}_{k}\cosh(n\xi_{k})=na_{n,c}\sinh(n\xi_{1}),\\ &\sinh(n\xi_{l})\sum_{k=1}^{l-1}\phi^{n}_{k}\frac{1}{\mathrm{e}^{n\xi_{k}}}+\left(\lambda_{l}-\frac{1}{2\mathrm{e}^{2n\xi_{l}}}\right)\phi^{n}_{l}-\frac{1}{\mathrm{e}^{n\xi_{l}}}\sum_{k=l+1}^{N}\phi^{n}_{k}\cosh(n\xi_{k})=na_{n,c}\sinh(n\xi_{l}),l=2,3,\ldots,N-1,\\ &\sinh(n\xi_{N})\sum_{k=1}^{N-1}\phi^{n}_{k}\frac{1}{\mathrm{e}^{n\xi_{k}}}+\left(\lambda_{N}-\frac{1}{2\mathrm{e}^{2n\xi_{N}}}\right)\phi^{n}_{N}=na_{n,c}\sinh(n\xi_{N}).\end{aligned}\right$.
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with $A$ being the multi-layered confocal ellipses given by (4.2)–(4.3). Let $H$ be given by (4.11). If $\mathbf{M}_{N,c}^{(n)}$ is invertible, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
[…]
@@
It follows from Theorem 4.1, (2.7) and (4.4) that the perturbed electric potential uHu-H in $A_{l}$, $l=0,1,\ldots,N$, can be given by
@@
Let the multi-layered confocal ellipses $A=\cup_{k=1}^{N}A_{k}$ be given by (4.2)–(4.3). Assume that the background electrical potential $H$ admits the following multipolar expansion:
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with $A$ being the multi-layered confocal ellipses given by (4.2)–(4.3). Let $H$ be given by (4.13). If $\mathbf{M}_{N,s}^{(n)}$ is invertible, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
[…]
@@
One can see that the problem (2.3) is solvable if and only if the corresponding $n$-order GPM $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ is invertible. If $\Re\sigma>0$, then $\nabla\cdot\sigma(x)\nabla$ is an elliptic PDO (Partial Differential Operator), which guarantees the well-posedness of (2.3) and hence the invertibility of $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ . On the other hand, if for some layers the materials are negative, the ellipticity of $\nabla\cdot\sigma(x)\nabla$ might be broken. Nevertheless, if for those negative materials, say $\sigma_{k}$, one has that $\Im\sigma_{k}\neq 0$, then the invertibility of $\mathbf{M}_{N,c}^{(n)}$ and $\mathbf{M}_{N,s}^{(n)}$ can also be guaranteed. This can be seen in our subsequent analysis.
@@ 5.1 Plasmon resonance
As mentioned above, plasmon resonance is usually associated with some eigenvalue problem generated by the PDE system. According to our earlier discussion in Remark 4.1, if the parameters $\sigma_{k}$, $k=1,2,\ldots,N$, are all positive real-valued, the elastostatic system the conductivity transmission problem (2.1)–(2.4) has only trivial solution if H=0{H}=0. We seek non-trivial solutions to (2.1)–(2.4) when the conductivity in (2.10) is allowed to be negative valued, i.e.,
@@
By using Theorems 4.2 and 4.4, we can oatain the exact matrix representation of the NP operator in multi-layered confocal ellipses.
@@
Suppose $u$ is the solution to the conductivity transmission problem (2.1)–(2.4) with the conductivity $\sigma$ given by (2.10) and AA being the multi-layered confocal ellipses given by (4.2)–(4.3).
@@
    If $\lambda$ is not the eigenvalue of $\mathbf{K}_{N,c}^{(n)}$, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
@@
    If $\lambda$ is not the eigenvalue of $\mathbf{K}_{N,s}^{(n)}$, then the problem (2.3) is uniquely solvable with the solution given by the following formula:
@@
where $\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ or $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, then the plasmon resonance mode occurs.
@@
$\left\{\begin{aligned} &\left(\lambda_{1}-\frac{1}{2\mathrm{e}^{2n\xi_{1}}}\right)\phi^{n}_{1}-\frac{1}{\mathrm{e}^{n\xi_{1}}}\sum_{k=2}^{N}\phi^{n}_{k}\cosh(n\xi_{k})=0,\\ &\sinh(n\xi_{l})\sum_{k=1}^{l-1}\phi^{n}_{k}\frac{1}{\mathrm{e}^{n\xi_{k}}}+\left(\lambda_{l}-\frac{1}{2\mathrm{e}^{2n\xi_{l}}}\right)\phi^{n}_{l}-\frac{1}{\mathrm{e}^{n\xi_{l}}}\sum_{k=l+1}^{N}\phi^{n}_{k}\cosh(n\xi_{k})=0,l=2,3,\ldots,N-1,\\ &\sinh(n\xi_{N})\sum_{k=1}^{N-1}\phi^{n}_{k}\frac{1}{\mathrm{e}^{n\xi_{k}}}+\left(\lambda_{N}-\frac{1}{2\mathrm{e}^{2n\xi_{N}}}\right)\phi^{n}_{N}=0.\end{aligned}\right$.
@@
where $\bm{\phi}=\left(\phi^{n}_{1},\phi^{n}_{2},\ldots,\phi^{n}_{N}\right)$. If $\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$, by letting $\delta\rightarrow 0$, one has $\left|-\lambda^{*}\mathbf{I}_{N}-\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}\right|=0$ and thus the homogeneous linear equation (5.7) has non-trivial solutions. In view of $u(x)=\sum_{k=1}^{N}\mathcal{S}_{\Gamma_{k}}[\phi_{k}](x)$ and (4.12), one immediately has the problem (5.6) has non-trivial solutions.
If $\lambda^{*}$ is an eigenvalue of $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, one also has the problem (5.6) has non-trivial solutions in a similar manner. The proof is complete. ∎
@@
[…]